if the mpc is 0.75 what is the size of the multiplier and the total change in real gdp demanded following a $10 billion increase in spending?
Answer #1

Answer

Multiplier =1/(1-MPC)

=1/(1-0.75)

=4

change in real GDP=intial change in spending * multiplier

=10*4

=$40 billion

the GDP demanded increases by $40 billion